A plane flies 480 miles with the wind and 320 miles against the wind in the same length of time. If the speed of the wind is 28mph, what is the speed of the plane in still air?
Answer by checkley75(3666)

USING THE FORMULA (DISTANCE=RATE*TIME) WE HAVE:
480=(R+28)T & T=480/(R+28)
320=(R-28)T & T=320/(R-28)
SEEING AS THE TIMES ARE THE SAME WE HAVE THE EQUATION:
480/(R+28)=320/(R-28) NOW WE CROSS MULTIPLY & GET
480(R-28)=320(R+28)
480R-13440=320R+8960
480R-320R=8960+13440
160R=22400
R=22400/160
R=140 MPH FOR THE PLANE SPEED IN STILL AIR.
PROOF
480/(140+28)=480/168=2.8571428 HOURS
320/(140-28)=320/112=2.8571428 HOURS